You are selling tickets for a musical at your local community collage. Student tickets cost $5 and general admission tickets cost $8. If you sell 500 tickets and collect $3475, how many student tickets and how many general admission.
Answer by ptaylor(2198)

Let x= number of student tickets sold
Then 500-x has to equal the number of general admission tickets sold
Now we know that the amount of $ collected from selling student tickets x($5) plus the amount of $ collected from selling general admission tickets (500-x)($8) must equal the total amount collected ($3475). Thus, our equation to solve is:
x($5)+(500-x)($8)=$3475 Simplifying, we get:
5x+4000-8x=3475 and
-3x=-4000+3475
-3x=-525
x=175 student tickets
500-x=500-175=325 general admission tickets
Hope this helps --------------ptaylor